OpenStudy (goldphenoix):

Mr.H teaches 5 classes. The mean number of students per class is 17.4. If he teaches one more class, the mean increases to 18.5. How many students are in the sixth class?

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

let's say x is the total of the five classes so you have \[\frac x5 = 17.4\] agree?

OpenStudy (goldphenoix):

Hmm. I'm not sure. o.o.

OpenStudy (anonymous):

it said mean number students per class is 17.4 so you get the total students of the five classes and divide it by 5. that's how you get the mean right?

OpenStudy (goldphenoix):

Oh yes. I agree. It's the way you say 'total of the five classes" that confused me.

OpenStudy (anonymous):

okay so now it's settled that \[\frac x5 = 17.4\] and x is the total number of students in the 5 classes so now get x

OpenStudy (goldphenoix):

17.4 * 5 = 87 x = 87 ?

OpenStudy (anonymous):

right. now before i proceed, i need to clarify something

OpenStudy (goldphenoix):

Sure, what is it?

OpenStudy (anonymous):

the mean increases is 18.5 this is ambiguous...is that mean increases TO? or mean increase is?

OpenStudy (goldphenoix):

to* My bad. >_>

OpenStudy (anonymous):

then the next equation would be \[\frac{87 + y}6 = 18.5\] where y is the total number of students in the sixth class

OpenStudy (anonymous):

you get why that is the equation right?

OpenStudy (anonymous):

^ that looks like the right equation to me

OpenStudy (goldphenoix):

Yes, because there is a total of 87 students in the 5 classes, and an 'y' amount of student in the 6th class. Divide that by the total amount of class, 6. And it would be equal to 18.5 because that is the new mean after the 6th class is added.

OpenStudy (anonymous):

right

OpenStudy (goldphenoix):

Let see, to solve this you multiply 6 by 18.5 to get 111. So it should look like: 87 + y = 111 Subtract 87 to get: y = 24 There is 24 students in the 6th class.

OpenStudy (goldphenoix):

Alright, thanks. Now it make sense. Just an extra more step to solve this.

OpenStudy (anonymous):

that's the reason for the "interesting" comment